# LLM-JEPA: Large Language Models Meet Joint Embedding Predictive Architectures

### v1

@@ 3 Conclusion and Future Work
We introduced an alternative training objective for LLMs leveraging JEPAs. Our formulation is an exact replicate of the JEPA objective extensively used in vision–but that hadn’t been adapted to language yet. Crucially, our proposed LLM-JEPA maintains the generative capabilities of LLMs while improving their abstract prompt representation as empirically validated across datasets and models. While our experiments mostly focus on finetuning, preliminary pretraining experiment are promising which we plan to scale and more thoroughly test in future work. Regarding the limitations of LLM-JEPA, the main current bottleneck is the 3-fold compute cost during training required to obtain the representations of the views. We plan to explore possible mitigation that would mask the self-attention matrix and allow for our LLM-JEPA loss to be evaluated within a single forward pass through the LLM.
@@ Appendix A Appendix
Table 2: Generated samples by model pretrained by paraphrase dataset. The pretrained model is not good at terminating sentence. prompt and generation
@@ A.1 Faster LoRA Convergence
Table˜3 demonstrates that LoRA fine-tuning with $\mathcal{L}_{\rm LLM-JEPA}$ loss not only achieves substantially higher accuracy than using $\mathcal{L}_{\rm LLM}$ alone, but also converges more quickly. Notably, at a LoRA rank of $512$, our method already reaches accuracy comparable to full fine-tuning, whereas LoRA with only $\mathcal{L}_{\rm LLM}$ still exhibits a clear performance gap.
@@
Table 4: Pretraining ++ fine-tuning Llama-3.2-1B-Instruct accuracy on pretraining dataset paraphrase and fine-tuning dataset rotten_tomatoes and yelp by Next Token Prediction ($\mathcal{L}_{\rm LLM}$) loss vs. $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method). Note that $\mathcal{L}_{\rm LLM-JEPA}$ is applied only at pretraining. We tune $lr_{pre}$ and $lr_{ft}$ by $\mathcal{L}_{\rm LLM}$, and stick to them in LLM-JEPA pretraining. We run pretraining 5 times, and for each pretrained model, we run fine-tuning 5 times. Average accuracy and standard deviation are reported. We also report $p$-value of paired, single-tailed $t$-Test.
@@
rotten_tomatoes $\mathcal{L}_{\rm LLM}$ $56.57\pm 1.66$ 7.38e47.38e-4 $lr_{pre}=8e-5,lr_{ft}=4e-5$
@@
A.2 Ablation Study on the Role of $\mathcal{L}_{\rm LLM}$
@@
We vary the ratio $\gamma/\lambda$ within $[0,1]$ while enforcing $\max(\gamma,\lambda)=1$ to maintain a constant learning rate. Table˜5 shows that $\mathcal{L}_{\rm LLM}$ remains essential for generative performance: when $\gamma=0$, the fine-tuned model produces only empty outputs. This indicates that the JEPA component primarily serves as a regularization term, complementing the generative loss.
@@
A.3 Hyperparameter Tuning for LLM-JEPA
Despite its strong accuracy gains, LLM-JEPA introduces two additional hyperparameters. As shown in fig.˜3, the optimal configuration may occur at any point in the grid $(\lambda,k)\in\{0.5,1.0,2.0,4.0\}\times\{0,1,2,3,4\}$, which imposes a significant cost for hyperparameter tuning. While we have not identified an efficient method to explore this space, we empirically observe that adjacent grid points often yield similar accuracy, suggesting the potential for a more efficient tuning algorithm.
@@ A.4 Additional Generation Examples
Table˜7 presents additional examples generated by fine-tuning Llama-3.2-1B-Instruct on the NL-RX-SYNTH dataset using $\mathcal{L}_{\rm LLM}$ and $\mathcal{L}_{\rm LLM-JEPA}$, respectively.
Ground Truth $\mathcal{L}_{\rm LLM}$ $\mathcal{L}_{\rm LLM-JEPA}$ (ours)
lines not having the string ’dog’ followed by a number, 3 or more times
 ((dog.*[0-9].*)3,)  ((dog.*[0-9].*)3,)  ((dog.*[0-9].*){3,})
lines containing ending with a vowel, zero or more times
lines with a number or a character before a vowel
lines with words with the string ’dog’, a letter, and a number
[…]
@@ A.5 Overfitting Behavior in LoRA Fine-Tuning
We also conducted experiments to examine whether LoRA fine-tuning with $\mathcal{L}_{\rm LLM\text{-}JEPA}$ exhibits similar resistance to overfitting. As shown in fig.˜5, accuracy under $\mathcal{L}_{\rm LLM\text{-}JEPA}$ generally continues to improve with additional epochs, whereas fine-tuning with $\mathcal{L}_{\rm LLM}$ shows clear signs of overfitting. Notably, the standard deviation is much higher than in full fine-tuning, likely reflecting the lower capacity of LoRA fine-tuning. An interesting pattern emerges: for $\mathcal{L}_{\rm LLM\text{-}JEPA}$, larger standard deviations often coincide with dips in accuracy, whereas for $\mathcal{L}_{\rm LLM}$ they tend to accompany accuracy spikes. This suggests that such fluctuations may be unreliable indicators of generalization quality.
@@ A.6 Structured Representations Induced by LLM-JEPA
We also examine the representation space to better understand how LLM-JEPA regularizes learned features. Specifically, we plot $t$-SNE embeddings for both $\operatorname{Text}$ and $\operatorname{Code}$ across three settings: the base model, a model fine-tuned with $\mathcal{L}_{\rm LLM}$, and a model fine-tuned with $\mathcal{L}_{\rm LLM\text{-}JEPA}$. As shown in fig.˜6, clear structure emerges after fine-tuning with $\mathcal{L}_{\rm LLM\text{-}JEPA}$. We hypothesize that $\mathcal{L}_{\rm LLM\text{-}JEPA}$ enforces structure in the representation space by constraining the mapping from $\operatorname{Enc}(\operatorname{Text})$ to $\operatorname{Enc}(\operatorname{Code})$ within a narrow subspace. If this is the case, the SVD decomposition of $\operatorname{Enc}(\operatorname{Text})-\operatorname{Enc}(\operatorname{Code})$ should yield significantly smaller singular values, which is confirmed in fig.˜7. Furthermore, we hypothesize that the mapping is approximately linear. To test this, we compute the least-squares regression error, and table˜10 supports this hypothesis. Together, these results suggest that LLM-JEPA promotes a near-linear transformation between $\operatorname{Text}$ and $\operatorname{Code}$ representations, which may underlie its accuracy improvements.
@@
(a) Base model: No fine-tuning
(b) Baseline: Fine-tuned by NTP loss
(c) LLM-JEPA (Ours) $k=0$
(d) LLM-JEPA (Ours) $k=1$
Figure 6: $t$-SNE plot of Text and Code representations in (a) Base mode without fine-tuning, (b) Baseline that is fine-tuned with NTP loss, (c) LLM-JEPA (ours) with $k=0$, and (d) LLM-JEPA (ours) with $k=1$. Clearly LLM-JEPA (ours) induced nice structure on the representations while fine-tuning with NTP loss disrupted the structure in the base model.
Figure 7: The top 100 singular values of $\operatorname{Enc}(\operatorname{Text})-\operatorname{Enc}(\operatorname{Code})$. The curves of LLM-JEPA (ours) are a few magnitudes lower than that of base model and regular fine-tuning, meaning the mapping from Text to Code are confined within a narrow subspace, fostering the nice structure we see in Figure 6
@@ A.7 Performance Across Model Sizes
We also evaluate LLM-JEPA across different model sizes. As shown in table˜11, we observe statistically significant improvements at all scales. Since there is no official 8B version of Llama-3.2, we instead use Llama-3.1-8B-Instruct, where performance collapsed due to the model’s difficulty in properly terminating regular expressions. To address this, we additionally evaluate using a startswith criterion—that is, a prediction is considered correct if the generated regular expression begins with the ground-truth expression, removing the need for exact termination. Under this metric, we again observe statistically significant accuracy improvements.

### v2

@@ 6 Conclusion and Future Work
We introduced an alternative training objective for LLMs leveraging JEPAs. Our formulation is an exact replicate of the JEPA objective extensively used in vision–but that hadn’t been adapted to language yet. Crucially, our proposed LLM-JEPA maintains the generative capabilities of LLMs while improving their abstract prompt representation as empirically validated across datasets and models. While our experiments mostly focus on finetuning, preliminary pretraining experiment are promising which we plan to scale and more thoroughly test in future work. Regarding the limitations of LLM-JEPA, the primary bottleneck at present is the 2-fold increase in compute cost during training, which is mitigated by random loss dropout.
Limitations Despite its strong accuracy gains, LLM-JEPA introduces two additional hyperparameters. As shown in fig.˜7, the optimal configuration may occur at any point in a grid $(\lambda,k)$, which imposes a significant cost for hyperparameter tuning. While we have not identified an efficient method to explore this space, we empirically observe that adjacent grid points often yield similar accuracy, suggesting the potential for a more efficient tuning algorithm.
@@ Appendix A Appendix
Table 7: Generated samples by model pretrained by cestwc/paraphrase dataset. The pretrained model is not good at terminating sentence. prompt and generation
@@ A.1 Faster LoRA Convergence
Table˜8 demonstrates that LoRA fine-tuning with $\mathcal{L}_{\rm LLM-JEPA}$ loss not only achieves substantially higher accuracy than using $\mathcal{L}_{\rm LLM}$ alone, but also converges more quickly. Notably, at a LoRA rank of $512$, our method already reaches accuracy comparable to full fine-tuning, whereas LoRA with only $\mathcal{L}_{\rm LLM}$ still exhibits a clear performance gap.
@@
Table 9: Pretraining ++ fine-tuning Llama-3.2-1B-Instruct accuracy on pretraining dataset cestwc/paraphrase and fine-tuning dataset Rotten Tomatoes and Yelp by Next Token Prediction ($\mathcal{L}_{\rm LLM}$) loss vs. $\mathcal{L}_{\rm LLM-JEPA}$ loss (our method). Note that $\mathcal{L}_{\rm LLM-JEPA}$ is applied only at pretraining. We tune $lr_{pre}$ and $lr_{ft}$ by $\mathcal{L}_{\rm LLM}$, and stick to them in LLM-JEPA pretraining. We run pretraining 5 times, and for each pretrained model, we run fine-tuning 5 times. Average accuracy and standard deviation are reported. We also report $p$-value of paired, single-tailed $t$-Test.
@@
Rotten Tomatoes $\mathcal{L}_{\rm LLM}$ $56.57\pm 1.66$ 7.38e47.38e-4 $lr_{pre}=8e-5,lr_{ft}=4e-5$
@@
A.2 LLM-JEPA Induces Structured Representation
We present additional $t$-SNE plots of $\rm Text$ and $\rm Code$ representations in fig.˜6, which show that different values of $k$ yield similar structural patterns. In contrast, standard fine-tuning appears to further disrupt the representation structure compared to the baseline model.
(a) Base model: No fine-tuning
(b) Baseline: Fine-tuned by NTP loss
(c) LLM-JEPA (Ours) $k=0$
(d) LLM-JEPA (Ours) $k=1$
Figure 6: $t$-SNE plot of ${\textnormal{f}}Text$ and $\rm Code$ representations in (a) Base mode without fine-tuning, (b) Baseline that is fine-tuned with NTP loss, (c) LLM-JEPA (ours) with $k=0$, and (d) LLM-JEPA (ours) with $k=1$. Clearly LLM-JEPA (ours) induced nice structure on the representations while fine-tuning with NTP loss disrupted the structure in the base model.
A.3 Ablation Study on the Role of $\mathcal{L}_{\rm LLM}$
@@
We vary the ratio $\gamma/\lambda$ within $[0,1]$ while enforcing $\max(\gamma,\lambda)=1$ to maintain a constant learning rate. Table˜10 shows that $\mathcal{L}_{\rm LLM}$ remains essential for generative performance: when $\gamma=0$, the fine-tuned model produces only empty outputs. This indicates that the JEPA component primarily serves as a regularization term, complementing the generative loss.
@@ A.4 Additional Generation Examples
Table˜11 presents additional examples generated by fine-tuning Llama-3.2-1B-Instruct on the NL-RX-SYNTH dataset using $\mathcal{L}_{\rm LLM}$ and $\mathcal{L}_{\rm LLM-JEPA}$, respectively.
@@ A.5 Overfitting Behavior in LoRA Fine-Tuning
We also conducted experiments to examine whether LoRA fine-tuning with $\mathcal{L}_{\rm LLM\text{-}JEPA}$ exhibits similar resistance to overfitting. As shown in fig.˜8, accuracy under $\mathcal{L}_{\rm LLM\text{-}JEPA}$ generally continues to improve with additional epochs, whereas fine-tuning with $\mathcal{L}_{\rm LLM}$ shows clear signs of overfitting. Notably, the standard deviation is much higher than in full fine-tuning, likely reflecting the lower capacity of LoRA fine-tuning. An interesting pattern emerges: for $\mathcal{L}_{\rm LLM\text{-}JEPA}$, larger standard deviations often coincide with dips in accuracy, whereas for $\mathcal{L}_{\rm LLM}$ they tend to accompany accuracy spikes. This suggests that such fluctuations may be unreliable indicators of generalization quality.
@@ A.6 Structured Representations Induced by LLM-JEPA
We also examine the representation space to better understand how LLM-JEPA regularizes learned features. Specifically, we plot $t$-SNE embeddings for both $\operatorname{Text}$ and $\operatorname{Code}$ across three settings: the base model, a model fine-tuned with $\mathcal{L}_{\rm LLM}$, and a model fine-tuned with $\mathcal{L}_{\rm LLM\text{-}JEPA}$. As shown in fig.˜4, clear structure emerges after fine-tuning with $\mathcal{L}_{\rm LLM\text{-}JEPA}$. We hypothesize that $\mathcal{L}_{\rm LLM\text{-}JEPA}$ enforces structure in the representation space by constraining the mapping from $\operatorname{Enc}(\operatorname{Text})$ to $\operatorname{Enc}(\operatorname{Code})$ within a narrow subspace. If this is the case, the SVD decomposition of $\operatorname{Enc}(\operatorname{Text})-\operatorname{Enc}(\operatorname{Code})$ should yield significantly smaller singular values, which is confirmed in fig.˜3. Furthermore, we hypothesize that the mapping is approximately linear. To test this, we compute the least-squares regression error, and table˜14 supports this hypothesis. Together, these results suggest that LLM-JEPA promotes a near-linear transformation between $\operatorname{Text}$ and $\operatorname{Code}$ representations, which may underlie its accuracy improvements.
@@ A.7 Performance Across Model Sizes
We also evaluate LLM-JEPA across different model sizes. As shown in table˜15, we observe statistically significant improvements at all scales. Since there is no official 8B version of Llama-3.2, we instead use Llama-3.1-8B-Instruct, where performance collapsed due to the model’s difficulty in properly terminating regular expressions. To address this, we additionally evaluate using a startswith criterion—that is, a prediction is considered correct if the generated regular expression begins with the ground-truth expression, removing the need for exact termination. Under this metric, we again observe statistically significant accuracy improvements.
@@
Figure 9: Fine-tuning HellaSwag with Llama-3.2-1B allows $\lambda$ to be scaled up to 1024, with performance continuing to improve.